# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ 3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (62). Moreover, under the perturbative regime of Proposition 1, one may take for example $C_{2}:=2\kappa+4\kappa^{2}M_{G}$.
Proof.
See Appendix D.3. ∎
@@
Figure 1 summarizes the overall two-stage design. Stage 2 is used purely as a warm-started post-processing map: it reuses the current stage 1 policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule. For numerical stability and amortized deployment, we additionally consider (i) a residual/control-variate form of the projection and (ii) interactive distillation of the projected rule into the policy network. These practical coupling mechanisms are described in Appendix F.
@@ 4 Breaking the Dimensional Barrier under Drift Uncertainty
Our goal is to test whether Pontryagin-guided learning and projection remain stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We compare Stage 1 (PG–DPO) to Stage 2 (Pontryagin projection), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
@@ 4.1 Benchmark market and evaluation protocol
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, decision-time evaluation uses $(t,X_{0})$ with fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (notably in Stage 2 projection). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
@@
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sqrt{D})\,dW_{t}^{\varepsilon},$ (66)
with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $D\in(0,\infty)^{d}$. This implies
$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(D)\;=\;FF^{\top}+\mathrm{diag}(D),$ (67)
@@
One-shot generation and fairness across methods. For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},D,\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants. Within a fixed $d$, we change only the uncertainty covariance $P$ (aligned vs. misaligned and the scale $s$ below). This isolates algorithmic effects from instance-to-instance randomness and makes the scaling comparisons controlled.
@@
$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(D),$ (70)
@@
where $\widetilde{B}$ is generated independently of $B$ (and can be orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(D)$. This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
@@
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$ for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $N_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $N_{\mathrm{MC}}=400\cdot d$ rollouts. Here $N_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1 for stochastic gradient estimates, and in Stage 2 for costate/mixed-moment estimators at queried states). Each trajectory draws its own latent $\theta\sim q$ at time 0 and its own Brownian path thereafter; we do not use a separate inner loop of $\theta$-samples per trajectory. All methods share the same discretization scheme (Euler) and the same market instances within each $d$; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
@@
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (31). We evaluate each method at $t=0$ with a fixed horizon $T=1.5$ on a fixed grid $\{X_{0}^{(i)}\}_{i=1}^{N_{\mathrm{eval}}}$ and report Euclidean RMSE to the analytic reference:
@@
$\mathrm{RMSE}(u_{0},\pi_{q,\gamma}^{\mathrm{const}}):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|u_{0}(X_{0}^{(i)},T)-\pi_{q,\gamma}^{\mathrm{const}}(T)\big\|^{2}\bigg)^{1/2},$ (71)
@@
where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $u_{0}(\cdot)$ denotes the decision-time action prescribed by the method (deployable $\theta$-blind output). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (68)–(70), the remaining subsections compare baseline Stage 1 PG–DPO, post-hoc Stage 2 P–PGDPO projection, and interactive distillation under matched simulation budgets.
@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Mixed-moment estimation and a decoupling approximation. A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{G}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional, and finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can lead to large Monte-Carlo variance once the subsequent linear solve is applied. For numerical stability and a uniform protocol across geometries, we therefore use a simple decoupling (independence) approximation for these mixed moments,
@@
(and similarly for other mixed products), which is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales. While this approximation is most valuable under misalignment—where direction mixing can amplify mixed-moment noise—it also performs well in aligned regimes (where mixed moments are typically easier to estimate), and in the CRRA benchmark below it does not alter the qualitative scaling conclusions: projection remains stable, and the observed misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups. [^1] We note, however, that in extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down, in which case one should revert to full mixed-moment estimation (possibly with larger budgets and/or regularized/certified projection).
Figure 2: Decision-time Euclidean RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window. Concretely, for each condition we take the median RMSE over the last evaluation snapshots recorded during training. Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare Stage 1 (deployable) and Stage 2 (post-hoc projection), with and without interactive distillation. Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) trajectory budgets.
@@
We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (31). We track (i) the Monte-Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (71) with fixed horizon $T=1.5$. Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the projection/teacher direction uses the $p_{\theta}$ aggregation (Section 3.2).
@@
What is compared in Figure 2. Stage 1 (PG–DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients. Stage 2 (P–PGDPO; Section 3.2) applies a $q$-aggregated Pontryagin projection to a Stage 1 checkpoint; we use the residual form of Section F.1. Interactive distillation (Section F.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable Stage 1 policy network.
@@
Thus Figure 2 separates projection quality (Stage 2: post-hoc projected, still $\theta$-blind) from amortized deployable quality (Stage 1 distilled: single forward pass).
@@
Figure 3: Pathwise sanity check at $d=100$ under common random numbers. Black: analytic Gaussian constant-fraction benchmark evaluated with the remaining horizon $\tau=T-t$ (fixed $q$). Blue: Stage 1 PG–DPO. Orange: on-the-fly Stage 2 P–PGDPO teacher (computed via Monte Carlo projection at visited states). Top/bottom: aligned/misaligned.
@@
Stage 2 projection versus amortization: scaling with dimension. For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (Stage 1 (distill.) below Stage 1) while leaving Stage 2 essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation reduces the policy-class approximation/optimization gap by injecting that signal into $\pi_{\varphi}$.
@@
In contrast, under the misaligned geometry the picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.

### v3

@@ 3.1 DPO via stochastic gradient ascent and a conditional BPTT–PMP correspondence
where one admissible explicit choice of $\delta_{\mathrm{BPTT}}$ is given in (56). Moreover, under the stable inversion regime above one may take, for example, $C_{2}:=2\kappa+2\kappa^{2}M_{g}.$
Remark 5 (Population vs. empirical residual; scope of the bound).
The residual $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ in (59) is defined using population mixed-moment objects $(A_{\pi^{\mathrm{warm}}},g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}})$. In practice we monitor the empirical analogue $\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{g}_{t}^{\mathrm{mix}}$ and enforce the stable inversion regime by ridge/diagnostics/skip (Appendix F.5). Relating $\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}$ to $\|\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{L^{2}(\mu)}$ amounts to controlling the estimator error in $(\widehat{A}_{t},\widehat{g}_{t}^{\mathrm{mix}})$ and the size of $\pi^{\mathrm{warm}}$ on $\mathcal{D}$; we do not pursue nonasymptotic concentration bounds here.
Proof sketch.
Define the population mixed-moment projection map on $\mathcal{D}$ by
$\mathcal{T}(\pi)\;:=\;-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}.$
Let the population projected policy be
$\pi^{\mathrm{proj}}\;:=\;\mathcal{T}(\pi^{\mathrm{warm}})\;=\;-A_{\pi^{\mathrm{warm}}}^{-1}g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}.$
Notation (for the proof only). To avoid clutter in this proof sketch, we write
$\widehat{A}:=\widehat{A}_{t},\hskip 18.49988pt\widehat{g}:=\widehat{g}_{t}^{\mathrm{mix}},$
as functions on the working domain $\mathcal{D}$ (with the dependence on $(t,x,y)$ suppressed), so that $\widehat{\pi}^{\mathrm{agg,mix}}=-\widehat{A}^{-1}\widehat{g}$ is exactly (53).
Write the stage-2 (stabilized) estimator-based projected policy as
$\widehat{\pi}^{\mathrm{agg,mix}}\;:=\;-\widehat{A}^{-1}\widehat{g}$,
[…]
  1. 1.
    Estimator-to-population stability (Appendix E.1). By the stated uniform stability/invertibility assumptions (in particular, $\widehat{A}$ is an $L^{\infty}(\mathcal{D})$-small perturbation of $A_{\pi^{\mathrm{warm}}}$ under the adopted stabilizers), Proposition 1 applies with $A=A_{\pi^{\mathrm{warm}}}$ and $g=g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$, yielding
    $\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\big\|_{L^{2}(\mu)}\;\leq\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$
    for some constant $C_{2}$ depending only on the uniform inverse bound $\kappa$ and uniform $L^{\infty}$ bounds on the population $g$ terms (one explicit admissible $\delta_{\mathrm{BPTT}}$ is given in (56)).
  2. 2.
    Slab-wise contraction of the population projection map (Appendix E.2). On each slab $\mathcal{D}_{k}$, the slab-wise Lipschitz gain assumption (57) together with the uniform bounds on $\mathcal{U}$ implies (via Proposition 2) that for all $\pi_{1},\pi_{2}\in\mathcal{U}$,
    $\|\mathcal{T}(\pi_{1})-\mathcal{T}(\pi_{2})\|_{k}\;\leq\;\varrho(\tau_{k})\,\|\pi_{1}-\pi_{2}\|_{k}\;\leq\;\varrho_{*}\,\|\pi_{1}-\pi_{2}\|_{k}.$
  3. 3.
    Fixed-point identity for the local interior optimum. Since $\pi^{\star,\mathrm{blind}}$ is a locally optimal interior deployable $\theta$-blind policy, it satisfies the aggregated stationarity condition on $\mathcal{D}$, equivalently the fixed-point relation
    $\pi^{\star,\mathrm{blind}}=\mathcal{T}(\pi^{\star,\mathrm{blind}}).$
    Therefore, for each slab $\mathcal{D}_{k}$,
    $\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;=\;\|\mathcal{T}(\pi^{\mathrm{warm}})-\mathcal{T}(\pi^{\star,\mathrm{blind}})\|_{k}\;\leq\;\varrho_{*}\,\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
  4. 4.
    Residual identity and small-gain closure. Let $r_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=A_{\pi^{\mathrm{warm}}}\pi^{\mathrm{warm}}+g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$ and $\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}:=\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\|_{k}$. By construction of $\pi^{\mathrm{proj}}$,
    $\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\;=\;A_{\pi^{\mathrm{warm}}}^{-1}\,r_{\mathrm{FOC,mix}}^{\mathrm{warm}},\hskip 18.49988pt\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}\;\leq\;\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$
    Using the triangle inequality and the contraction estimate above,
    $\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\|\pi^{\mathrm{warm}}-\pi^{\mathrm{proj}}\|_{k}+\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\kappa\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}+\varrho_{*}\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}.$
    Since $\varrho_{*}<1$, rearranging gives
    $\|\pi^{\mathrm{warm}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\frac{\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}},\hskip 18.49988pt\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{k}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm},k}^{\mathrm{mix}}.$
    Summing over slabs using $\|f\|_{L^{2}(\mu)}^{2}=\sum_{k}\|f\|_{k}^{2}$ yields
    $\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\|_{L^{2}(\mu)}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}.$
  5. 5.
    Conclusion. Finally, apply the triangle inequality:
    $\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\mathrm{proj}}\big\|_{L^{2}(\mu)}+\big\|\pi^{\mathrm{proj}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)},$
    and combine the bounds from Steps 1 and 4 to obtain (59).
A complete proof (including the precise stable-inversion conditions implied by the stage-2 stabilizers and an explicit admissible choice of $\delta_{\mathrm{BPTT}}$) is provided in Appendix E.3. ∎
@@
Figure 1 summarizes the overall two-stage PG–DPO design. Stage 2 is used purely as a warm-started post-processing map: it reuses the current Stage 1 (DPO) policy, estimates the adjoint blocks under frozen $\theta\sim q$, and outputs a single deployable $\theta$-blind projected rule. For numerical stability and amortized deployment, we additionally consider (i) a residual/control-variate form of the projection and (ii) interactive distillation of the projected rule into the policy network. These practical coupling mechanisms are described in Appendix G.
@@ 4 Breaking the Dimensional Barrier under Drift Uncertainty
Our goal is to test whether the two-stage PG–DPO pipeline remains stable as the number of assets grows. We generate APT-style covariance structures and sweep dimensions $d\in\{5,10,50,100\}$ under both aligned uncertainty ($P=s\Sigma$) and a misaligned geometry that rotates uncertainty away from market risk directions. We report results for (i) Stage 1 (DPO) warm-up policies, (ii) the Stage 2 Pontryagin projection (i.e., the post-processed output completing PG–DPO), and (when applicable) amortized variants via interactive distillation, under Monte Carlo budgets that scale linearly with $d$.
@@ 4.1 Benchmark market and evaluation protocol
$\theta$-blind deployability (and what uses $\theta$). Throughout Section 4, all reported policies are deployable and $\theta$-blind: the control is a function of observable state only (here, the observable decision-time state is $(t,x)$ with $t=0$ and a fixed horizon $T$), and never takes the realized latent premium $\theta$ as an input. The latent $\theta\sim q$ is sampled only inside the simulator to generate trajectories and to form Monte Carlo averages that approximate $q$-expectations (in Stage 1 for stochastic gradients, and in Stage 2 for aggregated adjoint/mixed-moment inputs). Any $\theta$-indexed objects (when referenced elsewhere) are used only for offline diagnostics and are not part of the deployable decision rule.
@@
$dR_{t}:=\frac{dS_{t}}{S_{t}}-r\mathbf{1}\,dt=\theta\,dt+B\,\Sigma_{f}^{1/2}\,dW_{t}^{f}+\mathrm{diag}(\sigma_{\varepsilon})\,dW_{t}^{\varepsilon},$ (60)
with $B\in\mathbb{R}^{d\times k_{\Sigma}}$, $\Sigma_{f}\succ 0$, and $\sigma_{\varepsilon}\in(0,\infty)^{d}$. This implies
$\Sigma=B\,\Sigma_{f}\,B^{\top}+\mathrm{diag}(\sigma_{\varepsilon}^{2})\;=\;FF^{\top}+\mathrm{diag}(\sigma_{\varepsilon}^{2}),$ (61)
@@
One-shot generation and fairness across methods. For each dimension $d$, we generate a single market instance $(B,\Sigma_{f},\sigma_{\varepsilon},\lambda_{m})$ once (using a fixed random seed) and hold it fixed across all algorithmic comparisons and MC-budget variants. Within a fixed $d$, we change only the uncertainty covariance $P$ (aligned vs. misaligned and the scale $s$ below). This isolates algorithmic effects from instance-to-instance randomness and makes the scaling comparisons controlled.
@@
$P=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}+s\,\mathrm{diag}(\sigma_{\varepsilon}^{2}),$ (64)
@@
where $\widetilde{B}$ is generated independently of $B$ (and can be orthogonalized against the span of $B$ to enforce large principal angles). The factor term is rescaled to match the aligned factor magnitude under the same $s$; concretely, writing $\Sigma_{\mathrm{fac}}:=B\Sigma_{f}B^{\top}$ and $\widetilde{\Sigma}_{\mathrm{fac}}:=\widetilde{B}\,\widetilde{\Sigma}_{f}\,\widetilde{B}^{\top}$, we rescale $\widetilde{\Sigma}_{\mathrm{fac}}$ so that $\mathrm{tr}(\widetilde{\Sigma}_{\mathrm{fac}})=\mathrm{tr}(s\,\Sigma_{\mathrm{fac}})$ while keeping the diagonal term fixed as $s\,\mathrm{diag}(\sigma_{\varepsilon}^{2})$. This geometry increases heterogeneity across $\theta\sim q$ and makes mixed-moment estimation and subsequent linear-algebra steps more fragile, providing a stringent scalability test.
@@
Experiment grid and simulation budgets. We vary the number of assets over $d\in\{5,10,50,100\}$ and sweep three uncertainty magnitudes $s\in\{10^{-3},10^{-2},10^{-1}\}$ for both aligned and misaligned geometries. To keep Monte Carlo noise comparable across dimensions, we use linear-in-$d$ trajectory budgets: a base regime with $M_{\mathrm{MC}}=100\cdot d$ rollouts and a high regime with $M_{\mathrm{MC}}=400\cdot d$ rollouts. Here $M_{\mathrm{MC}}$ denotes the number of independent simulated trajectories used to approximate the relevant $q$-expectations (in Stage 1/DPO for stochastic gradient estimates, and in Stage 2 for adjoint/mixed-moment estimators at queried states). Each trajectory draws its own latent $\theta\sim q$ at time 0 and its own Brownian path thereafter; we do not use a separate inner loop of $\theta$-samples per trajectory. All methods share the same discretization scheme (Euler) and the same market instances within each $d$; implementation details (network architecture, optimizer settings, and exact sampling conventions for Stage 1 vs. Stage 2) are reported in the implementation appendix and code release.
@@
Analytic reference and decision-time evaluation. In the static Gaussian benchmark, the analytic decision-time reference under constant portfolios is available in closed form. We use this closed-form rule only as an external decision-time target for evaluation; training does not impose the constant-portfolio restriction, and all methods learn from simulated trajectories over $[0,T]$ under the same $\theta$-blind constraint. For $\gamma>1$ we use the CRRA reference (30) at $(t,\tau)=(0,T)$, i.e., $\pi_{q,\gamma}^{\mathrm{const}}(0,T)$.
We evaluate each method at $t=0$ with a fixed horizon $T=1.5$ on a fixed grid $\{X_{0}^{(i)}\}_{i=1}^{N_{\mathrm{eval}}}$ and report Euclidean RMSE to the analytic reference:
@@
$\mathrm{RMSE}\!\big(\pi_{0}^{\mathrm{meth}},\pi_{q,\gamma}^{\mathrm{const}}\big):=\bigg(\frac{1}{N_{\mathrm{eval}}}\sum_{i=1}^{N_{\mathrm{eval}}}\big\|\pi_{0}^{\mathrm{meth}}\!\big(X_{0}^{(i)};T\big)-\pi_{q,\gamma}^{\mathrm{const}}(0,T)\big\|^{2}\bigg)^{1/2},$ (65)
@@
where $\|\cdot\|$ is the Euclidean norm on $\mathbb{R}^{d}$ and $\pi_{0}^{\mathrm{meth}}(\cdot;T)$ denotes the method’s deployable decision-time action at time 0 for the fixed horizon $T$ (a $\theta$-blind output rule). With the benchmark fixed and with $(m,\Sigma,P)$ constructed as in (62)–(64), the remaining subsections compare (i) the Stage 1 (DPO) warm-up policy, (ii) the Stage 2 Pontryagin projection output (completing PG–DPO), and (iii) interactive distillation variants, under matched simulation budgets.
@@ 4.2 High-dimensional CRRA benchmark: projection and amortization
Mixed-moment estimation and a decoupling approximation. A practical issue throughout our experiments (both aligned and misaligned) is the estimation of mixed moments across the latent parameter, such as $\mathbb{E}_{\theta\sim q}[p_{t}^{\theta}(z)\,\theta]$ (and analogous products entering $\widehat{g}_{t}^{\mathrm{mix}}$), because the costate $p_{t}^{\theta}(z)$ is $\theta$-dependent and high-dimensional. Finite-sample covariance between $p_{t}^{\theta}$ and $\theta$ can create large Monte Carlo variance, which is then amplified by the subsequent linear solve in the projection step. For numerical stability and a uniform protocol across geometries, we therefore use a simple decoupling (independence) approximation for these mixed moments,
@@
(and similarly for other mixed products). This approximation is exact when the relevant Pontryagin objects are effectively $\theta$-invariant and is accurate whenever $\mathrm{Cov}_{q}(p_{t}^{\theta}(z),\theta)$ is small relative to marginal scales. In this CRRA benchmark it does not change the qualitative scaling conclusions: projection remains stable in aligned regimes, and the misaligned degradation is consistent with residual growth and curvature mismatch rather than catastrophic mixed-moment blow-ups. [^1] In extreme uncertainty/misalignment—where $\theta$–costate dependence becomes pronounced—the decoupling can break down. In that case one should revert to full mixed-moment estimation (with larger budgets and/or regularized/certified projection).
Figure 2: Decision-time Euclidean RMSE at $t=0$ versus dimension $d$ (log scale), summarized by a tail median over the late-training window. Concretely, for each condition we take the median RMSE over the last evaluation snapshots recorded during training. Rows: uncertainty magnitude $s\in\{10^{-3},10^{-2},10^{-1}\}$. Columns: aligned vs. misaligned geometry. Curves compare the Stage 1 warm-up policy (DPO) to the Stage 2 output (Pontryagin projection; completing PG–DPO), with and without interactive distillation (amortization). Solid vs. dashed lines correspond to MC base ($100\cdot d$) vs. high ($400\cdot d$) trajectory budgets.
@@
We consider the CRRA benchmark with $\gamma=2$ under Gaussian drift uncertainty $q$ and evaluate against the analytic constant $q$-reference (30). We track (i) the Monte Carlo objective estimate $\widehat{J}$ during training and (ii) the decision-time error at $t=0$ via RMSE (65) with fixed horizon $T=1.5$. Because stochastic optimization produces non-monotone and noisy RMSE curves, we summarize each condition by a robust tail median: the median RMSE over the final evaluation snapshots in the late-training window. Unless stated otherwise, the Stage 2 projection / on-the-fly teacher uses the mixed-moment ($p$-weighted) aggregation of Section 3.2 (implemented with the decoupling approximation above).
@@
What is compared in Figure 2. Stage 1 (DPO; Section 3.1) trains a deployable $\theta$-blind policy $\pi_{\varphi}$ by maximizing $\widehat{J}$ via pathwise gradients through the Euler simulator. Stage 2 (Pontryagin projection; Section 3.2) post-processes a Stage 1 checkpoint by estimating (aggregated) adjoint inputs and applying the statewise linear solve; we use the residual form of Section G.1. The resulting projected control is the output of the full PG–DPO pipeline. Interactive distillation (Section G.2) treats the Stage 2 projected control as a teacher signal and amortizes it back into a deployable policy network.
@@
Thus Figure 2 separates projection quality (Stage 2 / PG–DPO: post-hoc projected, still $\theta$-blind) from amortized deployable quality (distilled Stage 1: single forward pass).
@@
Figure 3: Pathwise sanity check at $d=100$ under common random numbers. Black: analytic Gaussian constant-fraction benchmark evaluated with the remaining horizon $\tau=T-t$ (fixed $q$). Blue: Stage 1 (DPO) warm-up policy. Orange: on-the-fly Stage 2 teacher obtained by Pontryagin projection at visited states (PG–DPO). Top/bottom: aligned/misaligned.
@@
Stage 2 projection versus amortization: scaling with dimension. For the aligned geometry, small and moderate uncertainty ($s=10^{-3},10^{-2}$) yield a sharp Stage 2 reduction in decision-time error across all tested dimensions, bringing RMSE down to the $10^{-5}$$10^{-4}$ range, while Stage 1 (DPO) policies remain around $10^{-3}$. Interactive distillation consistently improves the deployable policy (distilled Stage 1 below basic Stage 1) while leaving the Stage 2 projection essentially unchanged, confirming the intended division of labor: Stage 2 supplies a structured stationarity-correction signal, and distillation injects that signal into $\pi_{\varphi}$ to reduce the remaining approximation/optimization gap.
@@
In contrast, under the misaligned geometry the picture becomes more heterogeneous. For small to moderate uncertainty ($s=10^{-3},10^{-2}$), Stage 2 still improves decision-time RMSE at small $d$, but its advantage shrinks with dimension and can approach the $10^{-3}$ level by $d=100$. For the largest uncertainty scale ($s=10^{-1}$), Stage 1 (DPO) becomes markedly less reliable, whereas Stage 2 remains substantially better, indicating that projection can act as a stabilizing correction even when end-to-end learning is stressed. Across settings, the base and high MC budgets tend to yield similar tail-median RMSE curves, suggesting that linear-in-$d$ scaling of simulation budgets is sufficient for stable comparisons in this benchmark.